# Learning to Synthesize Fault-Tolerant Quantum Circuits by Diagonalization

### v1

@@ II-D Synthesis as a Markov Decision Process
$Q^{\pi}(s_{t},a_{t})=r(s_{t},a_{t})+\gamma\mathbb{E}_{p}V^{\pi}(s_{t+1})$ (4)
which is the value of taking action $a_{t}$ when in state $s_{t}$ and then following policy $\pi$.
RL algorithms typically consist of two entities, an agent and an environment (illustrated in Figure 2). The agent is trained in order to maximize the cumulative rewards it receives from the environment. Learning the optimal value function leads to a synthesis policy which produces circuit implementations of unitaries that contain the fewest possible gates. For the scope of this paper, it is assumed that a neural network is used to learn either a value function (Equation 3) or a Q function (Equation 4). A policy can be obtained by simply taking the action which maximizes either function.
Table I compares various unitary synthesis approaches which can be used to target the Clifford+T gate set. Among search-based methods, the simulated annealing based Synthetiq tool empirically finds better implementations of unitaries than other methods [20]. Even so, Synthetiq fails to produce results for high precision implementations of complex unitaries as the space of circuits is too large to search. Synthesizing complex unitaries to high precision requires a mechanism for handling continuous rotations; this is possible in both analytical and diagonalization approaches.
Method Qubits Category Precision Non-random?
MCTS [25] 2-5 Search (RL) $10^{-1}$ No
Policy Iter. [24] 1 Search (RL) $10^{-2}$ No
DQN [23] 1-2 Search (RL) $10^{-3}$ No
Synthetiq [20] 1-4 Search (SA) $10^{-3}$ Yes
DVNU (ours) 1-3 Search (RL) $\leq 10^{-6}$ Diagonalizable
gridsynth [16] 1 Analytical - $R_{Z}$
QSD [14] 2+ Analytical - Yes
Table I: Unitary synthesis approaches for the Clifford+T gate set. Among RL approaches, DVNU handles more complex inputs and with diagonalization can synthesize to very high precision. Precision is Hilbert-Schmidt distance. Non-random indicates that unitaries from real benchmarks can be handled. Synthetiq uses Simulated Annealing and empirically outperforms other pure search-based tools at low precision. gridsynth optimally decomposes 1-qubit $R_{Z}$ unitaries to any precision. The QSD is an analytical method which can be used along with gridsynth to synthesize Clifford+T circuits.
III Diagonalization
Directly synthesizing a unitary by finding a sequence of gates which inverts it is intractable when high-precision is needed. This is because a large number of discrete gates are required to implement continuous rotations. To address this we shift the problem scope from inversion to diagonalization. This approach assumes there are continuous degrees of freedom in unitaries which cannot be handled by a search-based synthesis process. To fit the paradigm of search based synthesis as described in Section II-C, we first define the problem of diagonalization as an MDP. At time $t\in\{1,\dots,T\}$ in the synthesis process, the state $s_{t}\in S$ is defined by
$s_{t}=L_{t}U_{tar}^{\dagger}R_{t}$ (5)
where $L_{t},R_{t}$ are each sequences of discrete operations (as in Equation 1). The set of initial states contains the adjoint of target unitaries $S_{I}=\{U_{tar}^{\dagger}\}$. The set of terminal states is the set of all states which can be approximately inverted by a diagonal unitary matrix. This corresponds to the set of states satisfying
$d_{D}(s_{t})=\min_{\theta\in\mathbb{R}^{2^{n}}}d_{HS}\big{(}s_{t},\exp(-i\theta\cdot\hat{\sigma})\big{)}\leq\epsilon$ (6)
@@ III Diagonalization
$r(s_{t})=\begin{cases}+1&\text{if }s_{t}\in S_{T}\\ -1&\text{else}\end{cases}$ (7)
@@
$R_{T}s_{T}^{\dagger}L_{T}=R_{T}D_{\theta}^{-1}L_{T}.$ (8)
Figure 3 illustrates the general form of three-qubit circuits which satisfy this structure. As an example, the Toffoli gate can be diagonalized in just two Hadamard gates.
Figure 3: Diagonalization ansatz for three-qubit unitaries. The diagonal matrix corresponds to $12^{n}-$ 1 continuous degrees of freedom which can be efficiently implemented with $R_{Z}(\theta)$ and CNOT gates. The $L$ and $R$ subcircuits are made of discrete Clifford+T operations. A Toffoli only requires two H gates to be diagonalized and its diagonalization can be implemented with only CNOT and $R_{Z}(\pm\frac{\pi}{4})$ ($T$ or $T^{\dagger}$ gates).
This process can be be thought of as a Singular Value Decomposition where the left and right singular vector matrices are restricted to unitaries which can be efficiently implemented by searching through gate sequences from a discrete gate set (i.e. Clifford+T). Although our implementation of synthesis by diagonalization uses RL, other search-based synthesis methods are equally valid.
III-A Handling Diagonal Unitaries and $R_{Z}$ Gates
The diagonal operator $D_{\theta}$ captures continuous degrees of freedom that cannot be efficiently handled by the $L_{t}$ and $R_{t}$ unitaries. An ansatz implementing $D_{\theta}$ can be prepared using techniques from the synthesis diagonal unitaries and of Pauli strings [33, 34, 35]. These diagonal unitaries require only $CNOT$ and $R_{Z}$ gates. We assume that the $R_{Z}$ rotations will be implemented without ancilla in the Clifford+T gate set using gridsynth [16].
IV Reinforcement Learning with Unitaries
To diagonalize unitaries, we adopt a Reinforcement Learning approach called the Deep Value Network for Unitaries (DVNU). DVNU indirectly learns a diagonalization policy $\pi$ using Value Iteration [36]. Compared to the approaches described in [22, 23], our model-based method trains faster and achieves better synthesis performance.
We introduce advancements necessary for compiling more complex quantum programs. First, we introduce a global phase invariance operation. Invariance to global phase ensures that networks learn to associate values with a canonical form of a unitary state. Second, we introduce a unitary perturbation data augmentation. By applying small unitary disturbances to states during training, the network learns to associate values with regions centered around points on the Bloch sphere. The impact of these architectural changes are illustrated in Figure 5. An ablation study summarizes the effect on performance of all architectural changes (see Figure 7).
IV-A Modeling Synthesis
@@
IV-B A Global Phase Invariant Network Architecture
Two observables which differ only by a global phase cannot be distinguished experimentally [29]. When dealing with unitary operators numerically, we are forced to choose representations which may differ due to global phase. To ensure that equivalent unitaries are represented in the same way, we require global phase invariance.
Definition 2.
A function $g$ is said to be global phase invariant if for two unitaries $U,W\in\mathbb{U}(2^{n})$ such that $U=e^{i\theta}W$, it is the case that $g(U)=g(W)$.
We construct a new phase invariant neural network layer similar to the manner of [37], which showed how nonlinear activation functions may be made equivariant to complex phase. Figure 4 illustrates this process and a proof of its correctness can be found in the Appendix.
[…]
Figure 4: A unitary $U$ can be transformed to global phase invariant representation $W$ by 1)1)1 ) squaring $U$ element-wise, 2)2)2 ) computing and conjugating the normalized mean to obtain a unit magnitude complex number $\hat{\mu}^{*}$, 3)3)3 ) multiplying $\hat{\mu}^{*}\times U$, and 4)4)4 ) conditionally correcting the phase by an angle of $\pi$.
IV-C Unitary Perturbations for Robust Value Learning
@@
IV-D Neural Radiance Field Encoding
Unitary states are represented using $2\times 2^{n}\times 2^{n}$ real tensors, where $n$ is the number of qubits. As we focus only on cases of high-precision synthesis where $n\in\{1,2,3\}$, our input data is extremely information dense. Past work has shown that neural networks struggle to fit high frequency patterns in data [38, 39]. We adopt a positional encoding technique used widely in the field of novel view synthesis and introduced by the NeRF project [40]. Input values are transformed by a map $\gamma:\mathbb{R}\rightarrow\mathbb{R}^{2L}$ such that
$\begin{split}\gamma(x_{i})=\Big{(}\sin(2^{0}\pi x_{i}),\cos(2^{0}\pi x_{i}),\dots,\\ \sin(2^{L-1}\pi x_{i}),\cos(2^{L-1}\pi x_{i})\Big{)}\end{split}$ (9)
@@ IV-D Neural Radiance Field Encoding
IV-E Deep Value Network for Unitary Synthesis
Figure 5: Visualization of global phase invariance and unitary perturbative data augmentations. Left: Without global phase invariance, the Bloch sphere is rotated so that equivalent states have different representations. Center: With no unitary perturbations on the inputs, reachable states are only represented as single points. Right: Global phase invariance and unitary perturbations ensure that states are represented consistently and that the network learns to associate entire $\epsilon$-regions with values.
@@ IV-E Deep Value Network for Unitary Synthesis
We leverage deep value iteration to learn gate set specific value functions. We refer to this architecture as the Deep Value Network for Unitaries (DVNU). This contrasts with past work such as [22, 23] which leverages a Deep Q Network (DQN) [41]. To train a DQN agent, a state-action value (Equation 4) must be learned for each $(s_{t},a_{t})$ pair. To ensure adequate training, the agent must receive examples of many different actions taken in each state. When the number of actions (in this case the gate set $|A|$) is large, the sample inefficiency of the DQN algorithm limits training progress. In comparison, the DVNU must only learn the value of a state $s_{t}$. The policy used is
$\pi(a|s_{t})=\arg\max_{a\in A}V^{\pi}_{\theta}(a\times s_{t})$ (10)
@@
Value iteration [42] is possible because synthesis can be modeled exactly and the reward is a deterministic function of the state. The previous work of [21] also employed a method based on value iteration for topological quantum computing. More recently [24] demonstrated how in restricted cases tabular (as opposed to deep RL) policy iteration can be used to solve for optimal synthesis policies.
The DVNU additionally leverages global phase invariance (Sec IV-B) and unitary perturbative data augmentations during training (Sec IV-C). As with many of these methods, DVNUs can be extended to be topology aware: the agent will only propose gates that are allowed by a machine’s qubit connectivity. As the learned value function is gate set and topology dependent, a new agent must be trained in order to handle new topologies and new basis gates.
Figure 6: Accuracy of global phase invariant and non-invariant architectures as a function of phase shift. Two otherwise identical networks are trained to match unitaries with their associated Clifford+T gates. Global phase shifts (angles on the unit circle) are applied to unitaries during training and testing. The invariant architecture maintains perfect accuracy (keeps a radius of $100\%$); it correctly matches unitaries with their associated gate no matter the global phase.
V Experiments
@@
In this section we evaluate our Reinforcement Learning diagonalization approach to quantum circuit synthesis detailed in Section IV. We show how our method can synthesize high-precision unitaries which are out of reach for other synthesis approaches.
Figure 7: Ablation study for 1-qubit (left) and 3-qubit (right) synthesis. Lines are the mean of four runs, shaded regions depict extrema. The $y$-axis combines agent success and problem difficulty. Global phase invariance can be learned in the simplest cases, but is necessary for harder problems. Invariant architectures outperform non-invariant ones given enough training time.
V-A How does global phase invariance impact performance?
@@
To answer this question, we can compare the ability of invariant and non-invariant architectures to learn useful properties about unitaries. We take as a proxy task the problem of predicting which unitary is associated with which gate in a finite gate set. Two structurally identical multi-layer perceptron [43] networks are trained to assign gate labels to unitaries associated with the two-qubit Clifford+T gate set for 250 epochs. Figure 6 illustrates how the network which first applies the invariance operation detailed in Section IV-B is resilient to global phase disturbances; no matter the global phase shift it is able to correctly match unitaries with their associated gates.
@@
Global-phase invariance enables neural networks which operate on quantum datasets to achieve better performance. Actively addressing this symmetry removes the requirement that networks learn invariant representations. In the 1-qubit case, the impact of global-phase invariance is less evident; agents trained with and without dephasing achieve similar performance. The effect is more pronounced when tackling 3-qubit circuit synthesis (see Figure 7).
V-B How does DVNU compare to other RL approaches?
@@
Figure 8: Probability that a DQN agent (left) and a DVNU agent (right) successfully finds implementations of random 1-qubit Clifford+T unitaries by inversion. The Target Depth-axis shows the number of gates in 250 randomly generated compilation targets. Adding gates increases compilation difficulty. The Iteration Number-axis measures how much training time has passed. The vertical axis shows the probability that a compilation agent will successfully compile a target circuit. Our DVNU agent significantly outperforms state of the art deep learning based unitary synthesis work demonstrated in the literature.
@@
For this experiment, random unitary targets are obtained by randomly generating quantum circuits in the Clifford+T gate set consisting of $\{H,X,Y,Z,S,T,S^{\dagger},T^{\dagger}\}$ gates. Circuits are generated by randomly selecting a member of the 1-qubit Clifford group followed by either $T$ or $T^{\dagger}$. Random circuits are used during training because they can easily be generated and because the difficultly of synthesizing the unitaries associated with these circuits can be tuned by adding or removing gates. Agents in this experiment are trained to invert unitaries, meaning that the state representation used is $s_{t}=C_{t}U_{tar}^{\dagger}$.
@@
We train two models, a DQN as in [22] and a DVNU. Both networks were made invariant to global phase. Agents are trained for 150 iterations (about 6 hours) on a machine with an AMD EPYC 7763 CPU and an NVidia A100 GPU. Figure 8 illustrates the results of this experiment. After training, the 1-qubit DVNU agent was able to synthesize random Clifford+T circuits of depth 80 with a success rate of 54%. Comparisons to other RL synthesis projects are highlighted in Table I.
@@
V-C How well can the RL agent diagonalize unitaries?

### v2

@@ 2.3. Synthesis as a Markov Decision Process
(4) $Q^{\pi}(s_{t},a_{t})=r(s_{t},a_{t})+\gamma\mathbb{E}_{p}V^{\pi}(s_{t+1})$
which is the value of taking action $a_{t}$ when in state $s_{t}$ and then following policy $\pi$. Value functions differ from Q functions in that they only require a state, not a state-action pair.
RL algorithms typically consist of two entities, an agent and an environment (illustrated in Figure 3). The agent is trained in order to maximize the cumulative rewards it receives from the environment. Learning the optimal value function leads to a synthesis policy which produces circuit implementations of unitaries that contain the fewest possible gates. For the scope of this paper, a neural network is used to learn either a value function or a Q function. A policy can be obtained by simply taking the action which maximizes either function.
3. Diagonalization
Directly synthesizing a unitary by finding a sequence of gates which inverts it is intractable when high-precision is needed. This is because a large number of discrete gates are required to implement any single continuous rotation. To address this we shift the problem scope from inversion to diagonalization. This approach assumes there are continuous degrees of freedom in unitaries which cannot be handled by a search-based synthesis process. To fit the paradigm of search-based synthesis as described in Section 2.2, we first define the problem of diagonalization as an MDP. At time $t\in\{0,1,\dots,T\}$ in the synthesis process, the state $s_{t}\in S$ is defined by
(5) $s_{t}=L_{t}U_{tar}^{\dagger}R_{t}$
where $L_{t},R_{t}$ are each sequences of discrete operations (as in Equation 1). The set of initial states contains the adjoint of target unitaries $S_{I}=\{U_{tar}^{\dagger}\}$, meaning $R_{0}=L_{0}=I_{2^{n}}$. The set of terminal states is the set of all states which can be approximately inverted by a diagonal unitary matrix. This corresponds to the set of states satisfying
(6) $d_{D}(s_{t})=\min_{\theta\in\mathbb{R}^{2^{n}}}d_{HS}\big{(}s_{t},\exp(-i\theta\cdot\hat{\sigma})\big{)}\leq\epsilon$
@@ 3. Diagonalization
(7) $r(s_{t})=\begin{cases}+1&\text{if }s_{t}\in S_{T}\\ -1&\text{else}\end{cases}$
@@
(8) $R_{T}s_{T}^{\dagger}L_{T}=R_{T}D_{\theta}^{-1}L_{T}.$
Figure 4 illustrates the general form of three-qubit circuits which satisfy this structure. As an example, the Toffoli gate can be diagonalized by just two Hadamard gates.
Figure 4. Diagonalization ansatz for three-qubit unitaries. The diagonal matrix corresponds to $2^{n}$-$1$ continuous degrees of freedom which can be efficiently implemented with $R_{Z}(\theta)$ and CNOT gates. The $L$ and $R$ subcircuits are made of discrete Clifford+T operations. A Toffoli only requires two H gates to be diagonalized and its diagonalization can be implemented with only CNOT and $R_{Z}(\pm\frac{\pi}{4})$ ($T$ or $T^{\dagger}$ gates).
This process can be be thought of as a Singular Value Decomposition where the left and right singular vector matrices are restricted to unitaries which can be efficiently implemented by searching through gate sequences from a discrete gate set (i.e. Clifford+T). Although our implementation of synthesis by diagonalization uses RL, other search-based synthesis methods are equally valid. A tool such as Synthetiq can be retrofitted to leverage diagonalization.
3.1. Handling Diagonal Unitaries and Rotation Gates
The diagonal operator $D_{\theta}$ captures continuous degrees of freedom that cannot be efficiently handled by the $L_{t}$ and $R_{t}$ unitaries. An ansatz implementing $D_{\theta}$ can be prepared using specialized algorithms for synthesizing diagonal unitaries (Bullock and Markov, 2004; Li et al., 2022; Welch et al., 2014). These diagonal unitaries require only $CNOT$ and $R_{Z}$ gates. In this paper, $R_{Z}$ rotations are implemented without ancilla in the Clifford+T gate set using gridsynth (Ross and Selinger, 2016), but alternative techniques can also be used.
4. Reinforcement Learning with Unitaries
To diagonalize unitaries, we adopt a Reinforcement Learning approach called the Diagonalizing Value Network for Unitaries (DVNU). DVNU indirectly learns a diagonalization policy $\pi$ using Value Iteration (Bellman, 1957). Compared to the approaches described in (Moro et al., 2021; Chen et al., 2022), our model-based method trains faster and achieves better synthesis performance.
We introduce advancements necessary for compiling more complex quantum programs. First, we introduce a global phase invariance operation. Invariance to global phase ensures that networks learn to associate values with a canonical form of a unitary state. Second, we introduce a unitary perturbation data augmentation. By applying small unitary disturbances to states during training, the network learns to associate values with regions centered around points on the Bloch sphere. The impact of these architectural changes are illustrated in Figure 6. An ablation study summarizes the effect on performance of all architectural changes (see Figure 8). These improvements to the neural network architecture are evaluated only on discrete gate synthesis tasks, but are likely to boost performance in other problem domains such as parameterized ansatz search and seeded synthesis (Weiden et al., 2023).
4.1. Modeling Synthesis
@@
4.2. A Global Phase Invariant Network Architecture
Two observables which differ only by a global phase cannot be distinguished experimentally (Nielsen and Chuang, 2011). When dealing with unitary operators numerically, we are forced to choose representations which may differ due to global phase. To ensure that equivalent unitaries are represented in the same way, we require global phase invariance.
Definition 4.2.
A function $g$ is said to be global phase invariant if for two unitaries $U,W\in\mathbb{U}(2^{n})$ and $\theta\in\mathbb{R}$ such that $U=e^{i\theta}W$, it is the case that $g(U)=g(W)$.
We construct a new phase invariant neural network layer similar to the manner of (Singhal et al., 2021), which showed how nonlinear activation functions may be made equivariant to complex phase. Figure 5 illustrates this process and a proof of its correctness can be found in the Appendix.
Figure 5. A unitary $U$ can be transformed to global phase invariant representation $W$ by 1)1)1 ) squaring $U$ element-wise, 2)2)2 ) computing and conjugating the normalized mean to obtain a unit magnitude complex number $\hat{\mu}^{*}$, 3)3)3 ) multiplying $\hat{\mu}^{*}\times U$, and 4)4)4 ) conditionally correcting the phase by an angle of $\pi$. The Appendix contains further details.
4.3. Unitary Perturbations for Robust Value Learning
@@
Figure 6. Visualization of global phase invariance and unitary perturbative data augmentations. Points on the Bloch sphere represent pure single-qubit quantum mechanical state vectors (Nielsen and Chuang, 2011). Left: Without global phase invariance, the Bloch sphere is rotated so that equivalent quantum states have different representations. Center: With no unitary perturbations on the inputs, reachable quantum states are only represented as single points. Right: Global phase invariance and unitary perturbations ensure that quantum states are represented consistently and that the network learns to associate entire $\epsilon$-regions with values.
4.4. Neural Radiance Field Encoding
Unitary states are represented using $2\times 2^{n}\times 2^{n}$ real tensors, where $n$ is the number of qubits. As we focus only on cases of high-precision synthesis where $n\in\{1,2,3\}$, our input data is extremely information dense. Past work has shown that neural networks struggle to fit high frequency patterns in data (Rahaman et al., 2019; Xu, 2020). We adopt a positional encoding technique used widely in the field of novel view synthesis and introduced by the NeRF project (Mildenhall et al., 2020). Input values are transformed by a map $\gamma:\mathbb{R}\rightarrow\mathbb{R}^{2L}$ such that
(9) $\begin{split}\gamma(x_{i})=\Big{(}\sin(2^{0}\pi x_{i}),\cos(2^{0}\pi x_{i}),\dots,\sin(2^{L-1}\pi x_{i}),\cos(2^{L-1}\pi x_{i})\Big{)}\end{split}$
@@ 4.4. Neural Radiance Field Encoding
4.5. Diagonalizing Value Network for Unitary Synthesis
@@
We leverage deep value iteration to learn gate set specific value functions. We refer to this architecture as the Diagonalizing Value Network for Unitaries (DVNU). This contrasts with past work such as (Moro et al., 2021; Chen et al., 2022) which leverage Deep Q Networks (DQN) (Mnih et al., 2013). To train a DQN agent, a state-action value (Equation 4) must be learned for each $(s_{t},a_{t})$ pair. To ensure adequate training, the agent must receive examples of many different actions taken in each state. When the number of actions (in this case the gate set $|A|$) is large, the sample inefficiency of the DQN algorithm limits training progress. In comparison, the DVNU agent must only learn the value of a state $s_{t}$. The policy used is
(10) $\pi(a|s_{t})=\arg\max_{a\in A}V^{\pi}_{\theta}(a\times s_{t})$
@@
Value iteration (Sutton and Barto, 2018) is possible because synthesis can be modeled exactly and the reward is a deterministic function of the state. The previous work of (Zhang et al., 2020) also employed a method based on value iteration for topological quantum computing. More recently (Alam et al., 2023) demonstrated how in restricted cases, tabular (as opposed to deep RL) policy iteration can be used to solve for optimal synthesis policies.
5. Experiments
@@
In this section we evaluate our Reinforcement Learning diagonalization approach to quantum circuit synthesis detailed in Section 4. As inputs we consider randomly generated unitaries together with unitaries extracted from a collection of algorithms/programs used in many other synthesis studies. We show how our method can synthesize high-precision unitaries which are out of reach for other synthesis approaches.
A direct comparison of FT synthesis algorithms is complicated due to the trade-offs and intricacies of each method. At high levels of precision, most synthesis algorithms fail to find solutions. As a baseline, we consider the general approach where circuits are synthesized analytically using the Quantum Shannon Decomposition, followed by single-qubit FT synthesis of $R_{Z}$ gates. This is compared against our method applied to two and three qubit unitaries extracted from a number of benchmark circuits.
5.1. The Impact of Global Phase Invariance and Perturbations
@@
How do the proposed neural architectural modifications proposed in Section 4 impact performance? To answer this question, we can compare the ability of invariant and non-invariant architectures to learn useful properties about unitaries. We take as a proxy task the problem of predicting which unitary is associated with which gate in a finite gate set. Two structurally identical multi-layer perceptron (Rosenblatt, 1958) networks are trained to assign gate labels to unitaries associated with the two-qubit Clifford+T gate set for 250 epochs. Figure 7 illustrates how the network which first applies the invariance operation detailed in Section 4.2 is resilient to global phase disturbances; no matter the global phase shift it is able to correctly match unitaries with their associated gates.
@@
Global-phase invariance enables neural networks which operate on quantum datasets to achieve better performance. Actively addressing this symmetry removes the requirement that networks learn invariant representations. In the 1-qubit case, the impact of global-phase invariance is less evident; agents trained with and without dephasing achieve similar performance. The effect is more pronounced when tackling 3-qubit circuit synthesis (see Figure 8). The impact of the unitary perturbation data augmentation technique, and the advantage of using NeRF encodings, are both evident even when synthesizing single-qubit unitaries.
Figure 7. Accuracy of global phase invariant and non-invariant architectures as a function of phase shift. Two otherwise identical networks are trained to match unitaries with their associated Clifford+T gates. Global phase shifts (angles on the unit circle) are applied to unitaries during training and testing. The invariant architecture maintains perfect accuracy (keeps a radius of $100\%$); it correctly matches unitaries with their associated gate no matter the global phase.
Figure 8. Ablation study for 1-qubit (left) and 3-qubit (right) synthesis. Lines are the mean of four runs, shaded regions depict extrema. The $y$-axis scales agent success given a problem instance’s difficulty. Global phase invariance can be learned in the simplest cases (1-qubit synthesis), but is necessary for harder problems (2- and 3-qubit synthesis). Invariant architectures outperform non-invariant ones given enough training time.
5.2. Evaluating DVNU Against Other RL Approaches
@@
Figure 9. Probability that a DQN agent (left) and a DVNU agent (right) successfully finds implementations of random 1-qubit Clifford+T unitaries by inversion. The DQN learns a Q function (Eq 4) whereas the DVNU agent learns a value function (Eq 3). The Target Depth-axis shows the number of gates in 250 randomly generated compilation targets. Adding gates increases compilation difficulty. The Iteration Number-axis measures how much training time has passed. The vertical axis shows the probability that a compilation agent will successfully compile a target circuit. The architectural modifications made to the DVNU agent improve upon previously demonstrated state-of-the-art deep-learning-based unitary synthesis approaches.
@@
For this experiment, random unitary targets are obtained by randomly generating quantum circuits in the Clifford+T gate set consisting of $\{H,X,Y,Z,S,T,S^{\dagger},T^{\dagger}\}$ gates. Circuits are generated by randomly selecting a member of the 1-qubit Clifford group followed by either $T$ or $T^{\dagger}$. Random circuits are used during training because they can easily be generated and because the difficultly of synthesizing the unitaries associated with these circuits can be tuned by adding or removing gates. Agents in this experiment are trained to invert unitaries, meaning that the state representation used is $s_{t}=L_{t}U_{tar}^{\dagger}I_{2^{n}}$, and the distance function used is the Hilbert Schmidt distance (Eq 2).
@@
We train two models, a DQN (which learns a Q function) as in (Moro et al., 2021) and a DVNU (which learns a value function). Both networks were made invariant to global phase. Agents are trained for 150 iterations (about 6 hours) on a machine with an AMD EPYC 7763 CPU and an NVidia A100 GPU. Figure 9 illustrates the results of this experiment. After training, the 1-qubit DVNU agent was able to synthesize random Clifford+T circuits of depth 80 with a success rate of 54%. Comparisons to other RL synthesis projects are highlighted in Table 1.
@@
5.3. Comparing DVNU Synthesis to the Quantum Shannon Decomposition